In ΔABC ∠ACB = 120º and ∠CAB = 40º. AC is extended to P such that AP = AC + 2CB. The measure of ∠ABP is
 
1) 60º
2) 120º
3) 110º
4) 100º 

Grade:7

2 Answers

Arun
4 years ago
Here angle ACB=120 and angle CAB=40 degreeSo Angle ABC= 1801204020 degreeand using linear pair property, BCP=180120=60 degreeNow lets take mid point of CP as M and then CM=MP........(1)GivenAP=AC+2CBor AC+CP=AC+2CBCP=2CB........(2)CM+MP=2CBCM+CM=2CB2CM=2CBCM=CBmeans triangle CMB is isosceles, So in triangle BMC, CMB=CBMUsing angle sum property in  triangle BMCCMB+CBM=18060=120or CMB=CBM=60 degreeSo angle BMP= 18060=120 degree [Linear pair]So finally triangle BMCis equilateral triangle.So CM=MBNow using equation 1, MP=MBSo in triangle BMP, MBP=MPB.............(3)Using angle sum property in triangle MPB, MBP+MPB+120=180or MBP+MBP=60MBP=30 degreeSo angle ABP=20+60+30=110 degree
 
Rahul Sanapala
4 years ago
Seems like there is a problem. Your answer/solution is blank, could you please kindly repost the solution , Sir?